# Detecting ionized bubbles around luminous sources during the reionization era using HI 21-cm signal

### v1

@@ Abstract
Measuring the properties of the intergalactic medium (IGM) and sources during the Epoch of Reionization (EoR) is of immense importance. We explore the prospects of probing the IGM and sources through redshifted 21-cm observations of individual ionized bubbles surrounding known luminous sources during the EoR. Accordingly, we simulate HI 21-cm maps, foreground contaminants and system noise which are specific to the uGMRT and SKA1-Low observations. Following the subtraction of the foreground from the total visibility, we employ a visibility-based matched filter technique to optimally combine the desired HI 21-cm signal while minimizing the system noise. Our analysis suggests that these ionized bubbles can be detected with more than $5\sigma$ significance using approximately $\sim 2000$ and $\sim 3000$ hours of observation time with the uGMRT at redshift $7.1$ and $8.3$ respectively. The SKA1-Low should be able to detect these with more than $8\sigma$ significance using just $\sim 100$ hrs of observations. Further, we investigate the impact of the foreground subtraction on the detectability and find the signal-to-noise ratio decreases when smaller bandwidth is used. More importantly, we show that the matched filtering method can measure ionized bubble radius and constrain HI-neutral fraction reasonably well, helping deeper insights into source properties and the intergalactic medium.
@@ 1 Introduction
In recent times, there have been detections of several bright QSOs during the EoR [8, 9, 10, 11, 12, 13, 14, 15], and a handful of bright galaxies have also been detected at high redshifts as well [16, 17]. These bright, luminous sources are expected to produce large ionized bubbles (HII regions) around them within the IGM. At the initial and intermediate stages of reionization, isolated or nearly isolated ionized bubbles are expected, which will remain buried inside the neutral (or partially ionized) IGM. Observations of redshifted HI 21-cm signal are a promising tool to study the epoch of reionization, including the individual ionized regions [18, 19, 20, 21]. However, there are enormous challenges associated with these detections, as the signal is weaker by several orders of magnitude compared to the foreground contaminants [22, 23, 24, 25]. Moreover, the system noise is much stronger than the desired HI signal.
@@
Previous studies have explored the detection of HII regions in HI 21-cm images [26, 19, 27]. This method enables direct estimation of the HI neutral fraction, the sizes of ionized bubbles, and the properties of the underlying sources. However, direct imaging of HI using redshifted 21-cm signal during the Epoch of Reionization (EoR) is challenging with current instruments like the uGMRT, LOFAR, MWA due to their limited sensitivities. The SKA1-Low is anticipated to produce images of individual ionized bubbles using the HI 21-cm signal, but this requires considerably long observation times.
@@
This work investigates the prospects of individually detecting isolated ionized bubbles within the IGM, around known luminous sources using the matched filter algorithm. This technique works when the system noise of the radio interferometer dominates over the targeted HI signal. But the functional form of the targeted signal is known. The idea of applying a matched filter technique was first proposed in [28], which presents a visibility-based analytical framework to study the feasibility of detecting individual ionized bubbles during the EoR through radio-interferometric observations of redshifted HI 21-cm radiation. Later, it was shown that fluctuating intergalactic medium outside the targeted ionized bubble behaves as a noise and prohibits detection of ionized bubbles of small sizes with radius $\lesssim 6$ Mpc [29]. Subsequent work such as [30] used detailed numerical simulations to study the detection prospects and associated challenges. [31] presents scaling relations, which enable us to quickly estimate detection prospects for various ionized bubble sizes, redshifts, and instruments. Further, [32] used a Bayesian framework to explore the possibility of constraining the parameters that characterize the bubbles. This approach of detecting individual ionized bubbles in HI 21-cm maps is complementary to the widely followed approach of detecting the summary statistics of the signal such as the power spectrum, bispectrum, etc. While the statistical methods probe the large scale behavior of the reionizing sources and the intergalactic medium [33, 34, 35, 36, 37, 38, 39], our approach sheds light on individual sources, ionized regions, and surrounding neutral medium. Interpreting the detection is also straightforward in our approach, unlike the statistical methods which might be complicated and model-dependent.
@@
The outline of the paper is as follows: Section 2 provides a brief overview of the basics of the HI 21-cm signal around ionized bubbles. In section 3, we explain our methods to simulate the HI 21-cm signal around ionized bubbles, foreground contaminants, the generation of visibilities at baselines specific to a particular experiment, and system noise. Section 4 discusses foreground subtraction from the total visibilities and the application of the matched filter technique. Our results are presented in Section 5, while Section 6 presents a summary of the work undertaken in the paper alongside a discussion. Throughout our analysis we use the cosmological parameters $h=0.7$, $\Omega_{\rm m}=0.27$, $\Omega_{\Lambda}=0.73$, $\Omega_{\rm b}=0.044$ consistent with the WMAP measurements [40].
@@ 3.1 HI 21-cm signal around ionized bubbles
Here, we briefly describe simulated ionized regions surrounding a very bright quasar at $z=7.1$, and around a collection of galaxies at $z=8.3$ during the EoR. First, we simulate the dark-matter distribution using a $N$-body code  [46] within a cubic box of the comoving size of $\sim 527$ Mpc with $2048^{3}$ grids and the same number of particles. Then we use a Friends-of-Friends (FoF) [47] halo finder to identify the collapsed halos that host the galaxies/bright quasars. These galaxies/quasars serve as the sources of ionizing photons during the EoR. Subsequently, the resolution of simulated cubes is degraded to $512^{3}$ grids. This resolution resembles the corresponding angular resolution of the uGMRT and SKA1-Low and has the added benefit of reduced computational requirements for simulation. We then employ a semi-numerical prescription [48, 49, 50] on these outputs to simulate ionization maps and the HI 21-cm differential brightness temperature maps. We assume a flat cosmology model with $h=0.7$, $\Omega_{\rm m}=0.27$ and $\Omega_{\rm\Lambda}=0.73$ [40].
@@
The semi-numerical approach compares the average number of ionizing photons $\langle n_{\gamma}\rangle_{R}$ and neutral hydrogen atoms $\langle n_{\rm H}\rangle_{R}$ within a spherical region of radius $R$ around a grid cell. If the ionization condition $\langle n_{\gamma}\rangle_{R}\geq\langle n_{\rm H}\rangle_{R}$ is met for any radius $R$, then the grid cell is reckoned to be ionized, within that region. Otherwise, an ionized fraction value of $x_{\rm HII}=\langle n_{\gamma}\rangle_{\rm grid}/\langle n_{\rm H}\rangle_{\rm grid}$ is assigned to that grid cell. This procedure is repeated for varying radius $R$ up to a maximum value $R_{\rm max}$, over the whole simulation box. It is assumed that the number of ionizing photons from halos, $N_{\gamma}$ contributed to the ionization of the IGM follows the relation $N_{\gamma}\propto M_{\rm h}$, where $M_{\rm h}$ is the mass of the corresponding halo.
@@
Here, we consider two scenarios. First, we aim to simulate an ionized region around a very bright quasar at redshift $7.1$ similar to the one reported in [8]. In the second case, we simulate an ionized region around a collection of galaxies at redshift $8.3$ as reported in [17]. In the first scenario, we assume that such a very bright quasar is hosted by the most massive halo (of mass $M_{\rm h}\sim 6\times 10^{12},M_{\odot})$ found in our simulation [19]. We further assume that the ionizing photon emission rate by the QSO is $\sim 1.3\times 10^{57}\,{\rm s}^{-1}$ [30, 19]. We note that along with the central quasar, many galaxies reside in and outside the QSO ionized regions. These galaxies are also contributing ionizing photons. The mass-averaged neutral Hydrogen fraction resulting from galaxies outside the QSO HII region is $\langle x_{\rm HI}\rangle\sim 0.88$ at a redshift of $z=7.1$. In figure 1 (left panel), the resulting HI 21cm map is shown with the ionized bubble around the quasar in the center of the map. The ionized region around the central bright QSO is considerably bigger (of radius around $23.5$ comoving Mpc) than all other ionized regions produced by the galaxies.
@@
In the second case, we consider a collection of galaxies near the most massive halo (of mass $M_{\rm h}\sim 3.5\times 10^{12}M_{\odot}$) in our simulation at $z=8.3$. Simulation specifications such as the cubic box size, total number of particles used, and number of grids along a direction are the same as above. The ionizing luminosity of these galaxies, including the most massive halo in our simulation, is adjusted to create an HII bubble with a radius of $\sim 27.8$ Mpc, and the resulting mass-averaged neutral fraction is around $\langle x_{\rm HI}\rangle\sim 0.94$. The ionized bubble from this scenario is shown in the right panel of figure 1. We see that the ionized bubble in this scenario is slightly aspherical in shape, mainly because it is created by a collection of sources located at different locations inside the bubble.
@@
Figure 1: The left panel shows differential brightness temperature map of the HI 21-cm signal around an ionized bubble of comoving size $23.5$ Mpc at $z=7.1$ corresponding to observing frequency of $175$ MHz. The right panel shows similar map around an ionized bubble of comoving size $27.8$ Mpc at $z=8.3$ corresponding to observing frequency of $153$ MHz. The angular size of simulated maps is $\sim 3.3^{\circ}\times 3.3^{\circ}$ with $\sim 23^{\prime\prime}$ resolution.
@@ 3.2 Foreground simulations
The astrophysical foregrounds that contaminate the EoR HI 21-cm signal observations comprise various components [51], including diffuse galactic synchrotron emission (DGSE), radio synchrotron emission from point-like sources, free-free radio emission from diffuse ionized gas, etc. It has been shown that point sources dominate the $\sim 150$ MHz sky at the angular scales $\leq 4^{\circ}$ [25]. In our foreground simulation, we assume that point sources with fluxes exceeding a specific low threshold have been identified and removed from the visibility data. As a result, $150$ MHz and $175$ MHz skies are primarily dominated by the diffuse galactic synchrotron emission. We model the angular power spectrum corresponding to the diffuse galactic synchrotron emission using a power-law formulation [52] as,
@@
Figure 2: It shows a simulated 2D image of the diffuse galactic synchrotron emission (DGSE) of angular size $\sim 3.3^{\circ}\times 3.3^{\circ}$ with resolution of $\sim 23^{\prime\prime}$ at frequency $175$ MHz.
@@ 3.4 Simulating visibilities
We generate the mock visibility data of the HI 21-cm signal and the foregrounds at all simulated baselines at $256$ frequency channels, centered around frequencies $\nu=175$ MHz and $153$ MHz, using the Discrete Fourier Transformation (DFT) of the total simulated maps. We keep the baseline distribution the same in all frequency channels for simplicity.
@@ 3.5 System noise
Our mock simulation considers $8$ hrs of observations with $16$s and $320$s integration time for the uGMRT and SKA1-Low respectively. This produces a total of $783000$ and $11773440$ baselines for the uGMRT and SKA1-Low respectively. Lower integration time will produce a higher number of baselines for the SKA1-Low and hence will increase the requirement of computational resources which is beyond our reach. The system noise contribution for $8$ hrs of observations will be too high to detect the HI 21-cm signal around the individual ionized bubbles. Therefore, we present our results for $2048$ and $3072$ hrs for uGMRT observations at redshift $7.1$ and $8.3$ respectively, assuming a total of $256$ and $384$ nights of observations, with each night lasting $8$ hours. Assuming all nights are the same, we reduce the system noise corresponding to an $8$ hrs observation by a factor of $16$ and $19.6$ respectively for redshift $7.1$ and $8.3$ respectively. The resulting mock signal becomes equivalent to $2048$ and $3072$ hours of uGMRT observations. Similarly, for the SKA1-Low we present our results for $96$ hrs of observation for both redshift $7.1$ and $8.3$. Eventually, we obtain mock visibility data with all the signal contributions (HI 21-cm signal, galactic foreground, and system noise contribution) typical of the uGMRT and SKA1-Low observations.
@@
Figure 5 shows the total mock visibility as a function of frequency for two baselines $U=34\lambda$ (top left panel) and $U=230\lambda$ (top right panel) for the uGMRT at $175$ MHz. The corresponding mock visibility for the SKA1-Low is shown in the bottom panels (bottom left panel: $U=11\lambda$, and bottom right panel: $U=254\lambda$). They also show foreground contributions (red dashed lines). We see that the foreground dominates over the HI 21-cm signal and system noise at a small baseline (left panel) and, at a large baseline (right panel) the system noise dominates over the other two components. The HI 21-cm signal remains buried under the stronger foreground and system noise contributions. The total mock visibilities are similar for the second scenario at $z=8.3$ for both experiments, except for the fact that the system noise and foreground contributions are higher. Therefore, we do not show them here explicitly. The subsequent section discusses the subtraction of the foreground contribution.
@@
Figure 5: Upper panels show the total visibilities (HI signal + foregrounds + system noise) as a function of observing frequency for the uGMRT at baselines $U=34\lambda$ (upper left panel) and $U=230\lambda$ (upper right) at $175$ MHz. Lower panels show the same for the SKA1-Low at baselines $U=11\lambda$ (lower left panel) and $U=254\lambda$ (lower right). Red dashed lines show foreground contamination only.

### v2

@@ Abstract
Measuring the properties of the intergalactic medium (IGM) and sources during the Epoch of Reionization (EoR) is of immense importance. We explore the prospects of probing the IGM and sources through redshifted 21-cm observations of individual ionized bubbles surrounding known luminous sources during the EoR. Accordingly, we simulate HI 21-cm maps, foreground contaminants, and system noise which are specific to the uGMRT and SKA1-Low observations. Following the subtraction of the foreground from the total visibility, we employ a visibility-based matched filter technique to optimally combine the desired HI 21-cm signal while minimizing the system noise. Our analysis suggests that these ionized bubbles can be detected with more than $5\sigma$ significance using approximately $\sim 2000$ and $\sim 3000$ hours of observation time with the uGMRT at redshift $7.1$ and $8.3$, respectively, when the mean neutral hydrogen fraction outside the targeted bubble is $\bm{\sim 0.9}$. The SKA1-Low should be able to detect these bubbles with more than $8\sigma$ significance using only $\sim 100$ hrs of observations. The total observing time increases both for the uGMRT and SKA1-Low when the mean neutral hydrogen fraction outside the targeted bubble decreases. Further, we investigate the impact of foreground subtraction on the detectability and find the signal-to-noise ratio decreases when smaller bandwidth is used. More importantly, we show that the matched filtering method can measure ionized bubble radius and constrain HI-neutral fraction reasonably well, providing deeper insights into the source properties and the intergalactic medium.
@@ 1 Introduction
In recent times, there have been detections of several bright QSOs during the EoR [8, 9, 10, 11, 12, 13, 14, 15], and a handful of bright galaxies have also been detected at high redshifts as well [16, 17]. These bright, luminous sources will produce large ionized bubbles (HII regions) around them within the IGM. At the initial and intermediate stages of reionization, isolated or nearly isolated ionized bubbles are expected, which will remain buried inside the neutral (or partially ionized) IGM. Observations of redshifted HI 21-cm signal are a promising tool to study the epoch of reionization, including the individual ionized regions [18, 19, 20, 21, 22, 23, 24, 25, 26]. However, there are enormous challenges associated with these detections, as the signal is weaker by several orders of magnitude compared to the foreground contaminants [27, 28, 29, 30]. Moreover, the system noise is much stronger than the desired HI signal.
@@
Previous studies have explored the detection of HII regions in HI 21-cm images [31, 18, 32, 23, 24, 25, 26]. This method enables direct estimation of the HI neutral fraction, the sizes of ionized bubbles, and the properties of the underlying sources. However, direct imaging of HI using redshifted 21-cm signal during the Epoch of Reionization (EoR) is challenging with current instruments like the uGMRT, LOFAR, MWA due to their limited sensitivities. The SKA1-Low is anticipated to produce images of individual ionized bubbles using the HI 21-cm signal, but this requires considerably long observation times.
@@
This work investigates the prospects of individually detecting isolated ionized bubbles within the IGM, around known luminous sources using the matched filter algorithm. This technique works when the system noise of the radio interferometer dominates over the targeted HI signal. But the functional form of the targeted signal is known. The idea of applying a matched filter technique was first proposed in [33], which presents a visibility-based analytical framework to study the feasibility of detecting individual ionized bubbles during the EoR through radio-interferometric observations of redshifted HI 21-cm radiation. Later, it was shown that fluctuating intergalactic medium outside the targeted ionized bubble behaves as noise and hinders the detection of ionized bubbles of small sizes with radius $\lesssim 6$ Mpc [34]. Subsequent work such as [35] used detailed numerical simulations to study the detection prospects and associated challenges. [36] presents scaling relations, which enable us to quickly estimate detection prospects for various ionized bubble sizes, redshifts, and instruments. Further, [37] used a Bayesian framework to explore the possibility of constraining the parameters that characterize the bubbles. This approach of detecting individual ionized bubbles in HI 21-cm maps is complementary to the widely followed approach of detecting the summary statistics of the signal such as the power spectrum, bispectrum, etc. While the statistical methods probe the large scale behavior of the reionizing sources and the intergalactic medium [38, 39, 40, 41, 42, 43, 44], our approach sheds light on individual sources, ionized regions, and surrounding neutral medium. Interpreting the detection is also straightforward in our approach, unlike the statistical methods which might be complicated and model-dependent.
@@
The outline of the paper is as follows: Section 2 provides a brief overview of the basics of the HI 21-cm signal around ionized bubbles. In section 3, we explain our methods to simulate the HI 21-cm signal around ionized bubbles, foreground contaminants, the generation of visibilities at baselines specific to a particular experiment, and system noise. Section 4 discusses foreground subtraction from the total visibilities and the application of the matched filter technique. Our results are presented in Section 5, while Section 6 presents a summary of the work undertaken in the paper alongside a discussion. Throughout our analysis we use the cosmological parameters $h=0.7$, $\Omega_{\rm m}=0.27$, $\Omega_{\Lambda}=0.73$, $\Omega_{\rm b}=0.044$, $\bm{\sigma_{8}=0.83}$ and $\bm{n_{\rm s}=0.96}$ consistent with the WMAP measurements [45].
@@ 3.1 HI 21-cm signal around ionized bubbles
Here, we briefly describe simulated ionized regions surrounding a very bright quasar at $z=7.1$, and around a collection of galaxies at $z=8.3$ during the EoR. First, we simulated the dark-matter distribution using the $N$-body code [^2] https://github.com/rajeshmondal18/N-body developed by [51].  We simulated this within a cubic box of the comoving size of $\sim 527$ Mpc with $2048^{3}$ grids and the same number of particles. The minimum halo mass resolved in this simulation is $\bm{\sim 6.26\times 10^{9}M_{\odot}}$, corresponding to a minimum of $\bm{10}$ dark matter particles, each with a mass of $\bm{\sim 6.26\times 10^{8}M_{\odot}}$. Then we use a Friends-of-Friends  [^3] https://github.com/rajeshmondal18/FoF-Halo-finder (FoF) [52] halo finder to identify the collapsed halos that host the galaxies/bright quasars. These galaxies/quasars serve as the sources of ionizing photons during the EoR. Subsequently, the resolution of simulated cubes is degraded to $512^{3}$ grids. This resolution resembles the corresponding angular resolution of the uGMRT and SKA1-Low and has the added benefit of reduced computational requirements for simulation. We then employ a semi-numerical prescription [53, 54, 55] on these outputs to simulate ionization maps and the HI 21-cm differential brightness temperature maps.
@@
The semi-numerical approach compares the average number of ionizing photons $\langle n_{\gamma}\rangle_{R}$ and neutral hydrogen atoms $\langle n_{\rm H}\rangle_{R}$ within a spherical region of radius $R$ around a grid cell. If the ionization condition $\langle n_{\gamma}\rangle_{R}\geq\langle n_{\rm H}\rangle_{R}$ is met for any radius $R$, then the grid cell is reckoned to be ionized, within that region. Otherwise, an ionized fraction value of $x_{\rm HII}=\langle n_{\gamma}\rangle_{\rm grid}/\langle n_{\rm H}\rangle_{\rm grid}$ is assigned to that grid cell. This procedure is repeated for varying radius $R$ up to a maximum value $R_{\rm max}$, over the whole simulation box. It is assumed that the number of ionizing photons from halos, $N_{\gamma}$ contributed to the ionization of the IGM, follows the relation $N_{\gamma}\propto M_{\rm h}$, where $M_{\rm h}$ is the mass of the corresponding halo.
@@
Here, we consider two scenarios. First, we aim to simulate an ionized region around a very bright quasar at a redshift of $7.1$ similar to the one reported in [8]. In the second case, we simulate an ionized region around a collection of galaxies at redshift $8.3$ as reported in [17]. In the first scenario, we assume that such a very bright quasar is hosted by the most massive halo (of mass $M_{\rm h}\sim 6\times 10^{12},M_{\odot})$ found in our simulation [20]. We further assume that the ionizing photon emission rate by the QSO is $\sim 1.3\times 10^{57}\,{\rm s}^{-1}$ [35, 20]. We note that along with the central quasar, many galaxies reside in and outside the QSO ionized regions. These galaxies are also contributing ionizing photons. The mass-averaged neutral Hydrogen fraction resulting from galaxies outside the QSO HII region is $\langle x_{\rm HI}\rangle\sim 0.88$ at a redshift of $z=7.1$. In figure 1 (top left panel), the resulting HI 21cm map is shown with the ionized bubble around the quasar in the center of the map. The ionized region around the central bright QSO is considerably bigger (of radius around $23.5$ comoving Mpc) than all other ionized regions produced by the galaxies.
@@
In the second case, we consider a collection of galaxies near the most massive halo (of mass $M_{\rm h}\sim 3.5\times 10^{12}M_{\odot}$) in our simulation at $z=8.3$. Simulation specifications such as the cubic box size, total number of particles used, and number of grids along a direction are the same as above. The ionizing luminosity of these galaxies, including the most massive halo in our simulation, is adjusted to create an HII bubble with a radius of $\sim 27.8$ Mpc, and the resulting mass-averaged neutral fraction is around $\langle x_{\rm HI}\rangle\sim 0.94$. The ionized bubble from this scenario is shown in the top right panel of figure 1. We see that the ionized bubble in this scenario is slightly aspherical in shape, mainly because it is created by a collection of sources located at different locations inside the bubble.
In the above two cases, the targeted HII bubbles are isolated and nearly spherical. This geometry arises because the ionizing radiation from other sources is negligible, and the IGM outside the targeted bubbles remains highly neutral. While these scenarios may be suitable at the onset of reionization, they are likely unrealistic for the later stages of the process. At later stages of reionization, ionized bubbles around very bright sources are surrounded by neighboring ionized regions, leading to an aspherical shape for the targeted bubble. The third scenario accounts for this complexity. The bottom panel of Fig. 1 displays a 2D map of the differential brightness temperature ($\bm{\delta T_{b}}$) at redshift $\bm{z=7.1}$, where the central targeted HII bubble appears highly aspherical. The quasar at the center is the same as the first case. Surrounding it there are numerous other HII regions, unlike the first case. This results in a much lower mean neutral fraction, which is $\bm{0.52}$ in this scenario.
@@
Figure 1: The top left and bottom panels show differential brightness temperature maps of the HI 21-cm signal around an ionized bubble of comoving size $23.5$ Mpc at $z=7.1$ corresponding to observing frequency of $175$ MHz. The top right panel shows a similar map around an ionized bubble of comoving size $27.8$ Mpc at $z=8.3$ corresponding to observing frequency of $153$ MHz. The mass-averaged neutral hydrogen fraction resulting from the galaxies outside the central HII region is $\bm{0.88}$, $\bm{0.94}$, and, $\bm{0.52}$ for the top-left, top-right and bottom panels respectively. The angular size of simulated maps is $\sim 3.3^{\circ}\times 3.3^{\circ}$ with $\sim 23^{\prime\prime}$ resolution.
@@ 3.2 Foreground simulations
The astrophysical foregrounds that contaminate the EoR HI 21-cm signal observations comprise various components [56], including diffuse galactic synchrotron emission, radio synchrotron emission from point-like sources, free-free radio emission from diffuse ionized gas, etc.
The diffuse galactic synchrotron emission (DSGE) is one of the most dominant foreground components. We model the angular power spectrum corresponding to the diffuse galactic synchrotron emission using a power-law formulation [57] as,
@@
Figure 2: The left panel shows a simulated 2D image of the diffuse galactic synchrotron emission (DGSE) of angular size $\sim 3.3^{\circ}\times 3.3^{\circ}$ with resolution of $\sim 23^{\prime\prime}$ at frequency $175$ MHz. The right panel shows a 2D image of extragalactic radio point sources with a similar angular size, resolution and observing frequency.
3.2.1 Extragalactic radio point sources
One of the most significant foreground contaminants in HI 21-cm signal observation arises from the extragalactic radio point sources, such as Active Galactic Nuclei (AGN), quasars, and radio galaxies. At lower frequencies ($\bm{\sim 150}$ MHz), these sources dominate the sky at smaller angular scales $\bm{\lesssim 4^{\circ}}$ [30]. Therefore, in order to study the HII regions at frequencies of $\bm{150}$ and $\bm{175}$ MHz, it is essential to model the point sources efficiently.
In our simulations, we model the population of extragalactic radio point sources using the following $\bm{150}$ MHz differential source count model from [58]:
$\bm{\frac{dN}{dS}=\frac{10^{3.75}}{\rm Jy.Sr}\left(\frac{S}{1\,\rm Jy}\right)^{-1.6}},$ (3.3)
where $\bm{S}$ is the flux density of the point sources. We consider the point sources to be in the flux range between $\bm{0.1}$ mJy to $\bm{1}$ Jy and they are randomly distributed across an angular size of $\bm{\sim 3.3^{\circ}\times 3.3^{\circ}}$. Based on these parameters, we obtained $\bm{7775}$ point sources from the simulation. We model the spectral energy distribution of the point sources using a simple power law ($\bm{S\propto\nu^{-\alpha}}$), where the spectral index $\bm{\alpha}$ for each source is randomly assigned within the range of $\bm{0.7}$ to $\bm{0.8}$. The right panel of Figure 2 shows a 2D map of the point sources randomly distributed across an angular area of approximately $\bm{3.3^{\circ}\times 3.3^{\circ}}$, with an angular resolution of $\bm{\sim 23^{\prime\prime}}$.
@@ 3.4 Simulating visibilities
We generate the mock visibility data of the HI 21-cm signal and the foregrounds (DGSE and extragalactic point sources) at all simulated baselines at $256$ frequency channels, centered around frequencies $\nu=175$ MHz and $153$ MHz, using the Discrete Fourier Transformation (DFT) of the total simulated maps. We keep the baseline distribution the same in all frequency channels for simplicity.
@@ 3.5 System noise
Our mock simulation considers $8$ hrs of observations with $16$s and $320$s integration time for the uGMRT and SKA1-Low respectively. This produces a total of $783000$ and $11773440$ baselines for the uGMRT and SKA1-Low respectively. Lower integration time will produce a higher number of baselines for the SKA1-Low and hence will increase the requirement of computational resources which is beyond our reach. The system noise contribution for $8$ hrs of observations will be too high to detect the HI 21-cm signal around the individual ionized bubbles. Therefore, we present our results for $2048$ and $3072$ hrs of uGMRT observations at redshifts $7.1$ and $8.3$, respectively, assuming a total of $256$ and $384$ nights of observations, with each night lasting $8$ hours. Further, for the third scenario, we present the results at redshift $\bm{7.1}$ for $\bm{5120}$ hours, assuming $\bm{640}$ nights of observations. The increased observing time for the third scenario is due to the lower mean neutral hydrogen fraction outside the targeted bubble. Assuming all nights are the same, we reduce the system noise corresponding to an $8$ hrs observation by a factor of $16$, $19.6$, and $\bm{25.29}$ for the three scenarios, respectively. The resulting mock signal becomes equivalent to $2048$, $3072$, and $\bm{5120}$ hours of uGMRT observations. Similarly, for the SKA1-Low we present our results for $96$ hrs of observation for the first two cases at redshift $7.1$ and $8.3$ (neutral hydrogen fraction $\bm{\sim 0.9}$) and for $\bm{200}$ hrs of observation for the third case at redshift $\bm{7.1}$ (neutral hydrogen fraction $\bm{\sim 0.5}$). Eventually, we obtain mock visibility data with all the signal contributions (HI 21-cm signal, galactic foreground, extragalactic point sources, and system noise contribution) typical of the uGMRT and SKA1-Low observations.
@@
Figure 5 shows the mock visibility data comprising of HI signal, diffused galactic foreground contribution, and system noise as a function of frequency for two baselines $U=34\lambda$ (top left panel) and $U=230\lambda$ (top right panel) for the uGMRT at $175$ MHz. The corresponding mock visibility for the SKA1-Low is shown in the bottom panels (bottom left panel: $U=11\lambda$, and bottom right panel: $U=254\lambda$). They also show diffused galactic foreground contributions (red dashed lines). We see that the diffused galactic foreground dominates over the HI 21-cm signal and system noise at a small baseline (left panel) and, at a large baseline (right panel) the system noise dominates over the other two components. The HI 21-cm signal remains buried under the stronger foreground and system noise contributions.
@@
Figure 5: Upper panels show mock visibilities (HI 21-cm signal + diffuse galactic foreground + system noise) as a function of observing frequency for the uGMRT at baselines $U=34\lambda$ (upper left panel) and $U=230\lambda$ (upper right) at $175$ MHz. Lower panels show the same for the SKA1-Low at baselines $U=11\lambda$ (lower left panel) and $U=254\lambda$ (lower right). Red dashed lines show diffuse galactic foreground contamination only.
Figure 6 shows the total mock visibility comprising of HI signal, diffused galactic synchrotron emission, radiation from extra-galactic radio point sources, and system noise as a function of frequency for two baselines $\bm{U=34\lambda}$ (left panel) and $\bm{U=230\lambda}$ (right panel) for the uGMRT at $\bm{175}$ MHz. We note that Figure 6 shows contribution both from the extragalactic point sources, galactic diffuse synchrotron radiation along with other components unlike Fig. 5 which shows galactic diffuse synchrotron emission only as the foreground component. The radio point sources are modeled using a simplistic power-law form as discussed in subsection 3.2.1. The green lines indicate visibilities due to point sources. We see that at a small baseline (left panel) the diffuse galactic synchrotron emission dominates over the point sources; at a large baseline (right panel), the point sources dominate over the diffuse galactic synchrotron emission. The total mock visibilities are similar for the second scenario at $\bm{z=8.3}$ for both experiments, except for the fact that the system noise and foreground contributions are higher. Therefore, we do not show them here explicitly. The subsequent section discusses the subtraction of the foreground contribution.
Here, we would like to note that the visibility data used in this study does not contain calibration errors, ionospheric effects, and radio-frequency interference (RFI). A comprehensive investigation of these factors will be addressed in future work.
Figure 6: Left panel shows the total mock visibilities (HI 21-cm signal + DGSE + radio point sources + system noise) as a function of observing frequency for the uGMRT at baseline $\bm{U=34\lambda}$ and the right panel shows the same at baseline $\bm{U=230\lambda}$ at $\bm{175}$ MHz observing frequency. Red dashed lines show diffuse galactic foreground contamination and the green lines show the point source contribution separately.